An isomer of C5H12 gives total six isomeric product in chlorination . calc % yeild of the primary monochloro product which is chiral.REltivity


1 degreeC-H= 1


2 degree C-H=3


3 degree C-h=5

Grade:Upto college level

1 Answers

askiitians expert khajavali
14 years ago

probability factor = no. of hydrogens* reactivity
  
          in the case of chlorination the reactivities of  hydrogens are as follows
              1-degree c-h reactivity =1
             2-degree c-h reactivity =3.8
             3-degree c-h reactivity =5
already given 1degree s-h no.of hydrogens = 1 *1=1
                   2 degree c-h no of hydrogens = 3* 3.8=11.4
                    3-degree c-h no of hydrogens = 5*5 = 25

                      total     ==== 37.4
% yield of primary  mono chloro  pentane = no of primary hydrogens *100/37.4
                                                            = 2.6737 %
           secondary mono chloro pentane is chiral because it is have 4 different groups around the carbon

% yield of secondary mono chloro pentane  = 11.4*100/37.4
                                                               = 30%

% yield of tertiary  mono chloro pentane = 25*100/37.4
                                                         = 66.8%
as the % of the compound is more that is major product, more stable ,less reactine compound
 3. alkyl halide reacts with alkoxide to give ether it is a william son synthesis reaction